the width of a rectangle is 3feet less than the length. the perimeter is 62feet. find the length and width of the rectangle.
Answer by benazir.sj@gmail.com(109)

let the length of the rectangle =x
width =x-3
given ,
perimeter =62ft
2(l+w) =62ft
2(x+x-3) =62ft
2(2x-3)=62ft
4x-6 =62ft
4x=62+6 =68ft
x=68/4
x=17ft
therefore
length =17 ft
and
width =17-3
=14ft